# How Numerical Precision Affects Mathematical Reasoning Capabilities of LLMs

### v1

@@ C.1 Technical Lemmas for Logarithmic Precision MLP
We will first provide some basic results for logarithmic precision MLP, which comes from Feng et al. (2023). We omit the proofs here. The complete proof can be found in the appendix of Feng et al. (2023).
@@
For any ${\epsilon}>0$, there exists a two-layer MLP $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ with 4 hidden dimension and $\mathrm{GeLU}$ activation, such that for any $a,b\in[-M,M]$, $|f(a,b)-ab|\leq\epsilon$. Moreover, the $\ell_{\infty}$ norm of $f$ is bounded by $O(\mathrm{poly}(M,1/\epsilon))$.
@@
For any two-layer MLP ${\bm{g}}:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$ with $\mathrm{ReLU}$ activation, and the $l_{\infty}$ norm bounded by $M$. Then, for any $\epsilon>0$, there exists a two-layer MLP ${\bm{f}}$ with the same size and $\mathrm{GeLU}$ activation, such that for all ${\bm{x}}\in\mathbb{R}^{d_{1}}$, we have $\|{\bm{f}}({\bm{x}})-{\bm{g}}({\bm{x}})\|_{\infty}\leq\epsilon$. Moreover, the $\ell_{\infty}$ norm of ${\bm{f}}$ is bounded by $O(\mathrm{poly}(M,1/\epsilon))$.
@@
Consider selection function ${\bm{g}}:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d}$: ${\bm{g}}({\bm{x}},{\bm{y}},t)={\bm{x}}$ if $t>0$, and ${\bm{g}}({\bm{x}},{\bm{y}},t)={\bm{y}}$ otherwise. Then, for any $\epsilon>0$, $\alpha>0$, and $M>0$, there exist a two-layer MLP with $22d+$ 2 hidden dimension and $\mathrm{GeLU}$ activation, such that for all ${\bm{x}}\in[-M,M]^{d}$, ${\bm{y}}\in[-M,M]^{d}$, and $t\in(-\infty,-\alpha]\cup[\alpha,+\infty)$, $\|{\bm{f}}({\bm{x}},{\bm{y}},t)-{\bm{g}}({\bm{x}},{\bm{y}},t)\|_{\infty}\leq\epsilon$. Moreover, the $\ell_{\infty}$ norm is bounded by $O(\mathrm{poly}(M,1/\alpha,1/\epsilon))$.
@@ C.2 Technical Lemmas for Logarithmic Precision Attention Layer
Feng et al. (2023) also studied the expressive power of the standard attention layer and introduced two basic operations: COPY and MEAN, showing that the standard attention layer with log-precision is capable of these operations under some regularity assumptions. Here, we will provide the results and further consider a special operation called SINGLE COPY.
Consider a sequence of vectors ${\bm{x}}_{1},{\bm{x}}_{2},\cdots,{\bm{x}}_{n}$ where ${\bm{x}}_{i}=(\tilde{\bm{x}}_{i},r_{i},1)\in[-M,M]^{d+2}$ and $M$ is a constant. Denote attention matrices ${\bm{K}},{\bm{Q}},{\bm{V}}\in{\mathbb{R}}^{d^{\prime}\times(d+2)}$ and ${\bm{q}}_{i}={\bm{Q}}{\bm{x}}_{i}$, ${\bm{k}}_{j}={\bm{K}}{\bm{x}}_{j}$, ${\bm{v}}_{j}={\bm{V}}{\bm{x}}_{j}$. For any scalars $0<\rho,\delta<M$, define the matching set ${\mathcal{S}}_{i}=\{j\leq i:|{\bm{q}}_{i}\cdot{\bm{k}}_{j}|\leq\rho\}$ and consider the following operations:
  • COPY: The output is ${\bm{u}}_{1},\cdots,{\bm{u}}_{n}$ with ${\bm{u}}_{i}={\bm{v}}_{\mathrm{pos}(i)}$, where $\mathrm{pos}(i)=\operatorname{argmax}_{j\in{\mathcal{S}}_{i}}{r_{j}}$. The output ${\bm{u}}_{i}$ is undefined when $|{\mathcal{S}}_{i}|=0$.
  • MEAN: The output is ${\bm{u}}_{1},\cdots,{\bm{u}}_{n}$ with ${\bm{u}}_{i}=\operatorname{mean}_{j\in{\mathcal{S}}_{i}}{{\bm{v}}_{j}}$. The output ${\bm{u}}_{i}$ is undefined when $|{\mathcal{S}}_{i}|=0$.
  • SINGLE COPY: The output is ${\bm{u}}_{1},\cdots,{\bm{u}}_{n}$ with ${\bm{u}}_{i}={\bm{v}}_{\text{pos}(i)}$, where $\text{pos}(i)$ is the unique element in ${\mathcal{S}}_{i}$. The output is undefined if $|{\mathcal{S}}_{i}|\neq 1$.
We next make the following regularity assumption:
Assumption C.4.
For any input sequences ${\bm{x}}_{1},{\bm{x}}_{2},\cdots,{\bm{x}}_{n}$, the matrices ${\bm{Q}},{\bm{K}},{\bm{V}}$ and scalars $\rho,\delta$ satisfy:
  • For any $i,j\in[n]$, $|{\bm{q}}_{i}\cdot{\bm{k}}_{j}|\leq\rho$ or ${\bm{q}}_{i}\cdot{\bm{k}}_{j}\leq-\delta$ hold.
  • For any $i,j\in[n]$, $i=j$ or $|r_{i}-r_{j}|\geq\delta$ hold.
  • $\|{\bm{V}}\|_{\infty}\leq 1$.
Now, we will show a logarithmic precision attention layer with $O(d)$ embedding dimension and one attention head can perform the operations defined above.
@@
Suppose Assumption C.4 holds and $\rho\leq\frac{\delta^{2}}{8M}$. For any $\epsilon>0$, there exists an attention layer with a single attention head and $O(d)$ embedding dimension that can approximate $\mathrm{COPY}$ operation, and the $\ell_{\infty}$ norm of the parameters is bounded by $O(\mathrm{poly}(M,1/\delta,\log(n),\log(1/\epsilon)))$. Formally, for any input sequences ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, denote the attention output as ${\bm{o}}_{1},{\bm{o}}_{2},\dots,{\bm{o}}_{n}$. Then for any $i\in[n]$ with ${\mathcal{S}}_{i}\neq\varnothing$, we have $\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}\leq\epsilon$.
@@
Suppose Assumption C.4 holds and $\rho\leq\frac{\delta\epsilon}{16M\ln(\frac{4Mn}{\epsilon})}$. For any $0<\epsilon\leq M$, there exists an attention layer with a single attention head and $O(d)$ embedding dimension that can approximate $\mathrm{MEAN}$ operation, and the $\ell_{\infty}$ norm of the parameters is bounded by $O(\mathrm{poly}(M,1/\delta,\log(n),\log(1/\epsilon)))$. Formally, for any input sequences ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, denote the attention output as ${\bm{o}}_{1},{\bm{o}}_{2},\dots,{\bm{o}}_{n}$. Then for any $i\in[n]$ with ${\mathcal{S}}_{i}\neq\varnothing$, we have $\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}\leq\epsilon$.
We omit the proofs of Lemmas C.5 and C.6 here. The complete proof can be found in the appendix of Feng et al. (2023).
@@
Assume Assumption C.4 holds and $\delta-\rho\geq c\rho$ where $c>0$. For any $\epsilon>0$, there exists an attention layer with a single attention head and $O(d)$ embedding dimension that can approximate $\mathrm{SINGLE}$ $\mathrm{COPY}$ operation, and the $\ell_{\infty}$ norm of the parameters is bounded by $O(\mathrm{poly}(M,1/\delta,1/c,\log(n),\log(1/\epsilon)))$. Formally, for any input sequences ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, denote the attention output as ${\bm{o}}_{1},{\bm{o}}_{2},\dots,{\bm{o}}_{n}$. Then, for any $i\in[n]$ with $|{\mathcal{S}}_{i}|=1$, we have $\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}\leq\epsilon$.
@@
We construct the query, key and value as follows:
@@
where $\lambda>0$ is a constant. Denote $a_{ij}$ as the attention score, we have
$a_{i,j}=\frac{\exp(\lambda({\bm{q}}_{i}\cdot{\bm{k}}_{j}))}{\sum_{j^{\prime}}\exp(\lambda({\bm{q}}_{i}\cdot{\bm{k}}_{j^{\prime}}))}=\frac{\exp(\lambda({\bm{q}}_{i}\cdot{\bm{k}}_{j}))}{\sum_{j^{\prime}}\exp(\lambda({\bm{q}}_{i}\cdot{\bm{k}}_{j^{\prime}}))}.$
Since $\delta-\rho\geq c\rho$, we have $\delta-\rho\geq\frac{c}{c+1}\delta$. By setting $\lambda=\frac{(c+1)\ln(\frac{2nM}{\epsilon})}{c\delta}$ (which is bounded by $O(\mathrm{poly}(M,1/\delta,1/c,\log(n),\log(1/\epsilon)))$), we have
@@
Here, we use Assumption C.4 and $|{\mathcal{S}}_{i}|=1$ in (1), which implies that for $j^{\prime}\neq\mathrm{pos}(i)$, we have ${\bm{q}}_{i}\cdot{\bm{k}}_{j^{\prime}}\leq-\delta$; and in (2) we apply the inequality $\frac{1}{1+x}\geq 1-x$ for all $x\geq 0$. Thus we have
$\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}=\left\|\sum_{j}a_{ij}{\bm{v}}_{j}-{\bm{v}}_{\mathrm{pos}(i)}\right\|_{\infty}$ $\leq M\|{\bm{V}}\|_{\infty}\cdot\left(1-a_{i,\mathrm{pos}(i)}+\sum_{j\neq\mathrm{pos}(i)}a_{i,j}\right)$
$=M\|{\bm{V}}\|_{\infty}(2-2a_{i,\mathrm{pos}(i)})\leq\epsilon,$
which concludes the proof. ∎
C.3 Technical Lemmas for Constant Precision Calculation
In this section, we provide some technical lemmas for constant precision calculation and constant precision MLPs. Suppose we have $2s$-bit precision bits and no exponent bits, and denote $B_{s}=2^{s}-2^{-s}$. The maximum number we could represent is $B_{s}$, while the minimum is $-B_{s}$.
Lemma C.8 (Li et al., 2024, Lemma E.1 and E.2).
For any $s\in{\mathbb{N}}_{+}$, $\exp(-B_{s})=0$ and $\exp(B_{s})=B_{s}$.
@@ C.3 Technical Lemmas for Constant Precision Calculation
Notice that $\exp(B_{s})\geq\mathrm{e}B_{s}>2^{s+1}$, thus $\exp(-B_{s})\leq 2^{-s-1}$. This implies that $\exp(-B_{s})=0$. On the other hand, $\exp(B_{s})\geq B_{s}+1>B_{s}$, which implies that $\exp(B_{s})=B_{s}$. ∎
@@
For any $s\in{\mathbb{N}}_{+}$ , $\mathrm{GeLU}(-B_{s})=0$.
@@
It suffices to prove $B_{s}\Phi(-B_{s})\leq 2^{-s-1}$ where $\Phi$ is the CDF of Gaussian distribution.
For $s=1$, we have $B_{s}=\frac{3}{2}$, thus
@@
For $s\geq 2$, we have
$B_{s}\Phi(-B_{s})$ $=\frac{B_{s}}{\sqrt{2\pi}}\int_{B_{s}}^{+\infty}\mathrm{e}^{-\frac{x^{2}}{2}}\mathrm{d}x\leq\frac{B_{s}}{\sqrt{2\pi}}\int_{B_{s}}^{+\infty}\mathrm{e}^{-\frac{B_{s}x}{2}}\mathrm{d}x$
$\leq\sqrt{\frac{2}{\pi}}\mathrm{e}^{-\frac{B_{s}^{2}}{2}}\leq\frac{2\sqrt{2}}{\sqrt{\pi}(B_{s}^{2}+2)}\leq\frac{2\sqrt{2}}{\sqrt{\pi}2^{2s}}\leq\frac{1}{2^{s+1}},$
completing our proof. ∎

### v2

@@ C.1 Technical Lemmas for Logarithmic Precision MLP
In this subsection, we present several foundational results concerning logarithmic precision multi-layer perceptrons (MLPs), as introduced in Feng et al. (2023). For brevity, proofs of these results are omitted here but are available in the appendix of Feng et al. (2023).
@@
Let $\epsilon>0$. There exists a two-layer MLP $f:\mathbb{R}^{2}\to\mathbb{R}$ with four hidden units and $\mathrm{GeLU}$ activation, such that for any $a,b\in[-M,M]$, the inequality $|f(a,b)-ab|\leq\epsilon$ holds. Furthermore, the $\ell_{\infty}$ norm of $f$ is bounded by $O(\mathrm{poly}(M,1/\epsilon))$.
@@
Let ${\bm{g}}:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$ be a two-layer MLP with $\mathrm{ReLU}$ activation and $\ell_{\infty}$ norm bounded by $M$. Then, for any $\epsilon>0$, there exists a two-layer MLP ${\bm{f}}$ of the same size with $\mathrm{GeLU}$ activation such that for all ${\bm{x}}\in\mathbb{R}^{d_{1}}$, the inequality $\|{\bm{f}}({\bm{x}})-{\bm{g}}({\bm{x}})\|_{\infty}\leq\epsilon$ is satisfied. Moreover, the $\ell_{\infty}$ norm of ${\bm{f}}$ is bounded by $O(\mathrm{poly}(M,1/\epsilon))$.
@@
Consider the selection function ${\bm{g}}:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d}$ defined as
${\bm{g}}({\bm{x}},{\bm{y}},t)=\begin{cases}{\bm{x}}&\text{if }t>0,\\ {\bm{y}}&\text{otherwise}.\end{cases}$
For any $\epsilon>0$, $\alpha>0$, and $M>0$, there exists a two-layer MLP ${\bm{f}}$ with $22d+$ 2 hidden units and $\mathrm{GeLU}$ activation such that, for all ${\bm{x}}\in[-M,M]^{d}$, ${\bm{y}}\in[-M,M]^{d}$, and $t\in(-\infty,-\alpha]\cup[\alpha,+\infty)$, the inequality $\|{\bm{f}}({\bm{x}},{\bm{y}},t)-{\bm{g}}({\bm{x}},{\bm{y}},t)\|_{\infty}\leq\epsilon$ holds. Furthermore, the $\ell_{\infty}$ norm of ${\bm{f}}$ is bounded by $O(\mathrm{poly}(M,1/\alpha,1/\epsilon))$.
@@ C.2 Technical Lemmas for Logarithmic Precision Attention Layer
Feng et al. (2023) investigated the expressive power of the standard attention layer and introduced two fundamental operations: COPY and MEAN, demonstrating that a standard attention layer with logarithmic precision can perform these operations under certain regularity conditions. In this subsection, we restate their results and extend the discussion to a specialized operation referred to as SINGLE COPY.
Consider a sequence of vectors ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, where each ${\bm{x}}_{i}=(\tilde{{\bm{x}}}_{i},r_{i},1)\in[-M,M]^{d+2}$, and $M$ is a fixed constant. Let the attention matrices be ${\bm{K}},{\bm{Q}},{\bm{V}}\in\mathbb{R}^{d^{\prime}\times(d+2)}$, and define the following transformed vectors:
${\bm{q}}_{i}={\bm{Q}}{\bm{x}}_{i},\quad{\bm{k}}_{j}={\bm{K}}{\bm{x}}_{j},\quad{\bm{v}}_{j}={\bm{V}}{\bm{x}}_{j}.$
For any scalars $0<\rho,\delta<M$, define the matching set as:
${\mathcal{S}}_{i}=\{j\leq i:|{\bm{q}}_{i}\cdot{\bm{k}}_{j}|\leq\rho\}.$
Using this matching set, we define the following operations:
  • COPY: The output is a sequence of vectors ${\bm{u}}_{1},\dots,{\bm{u}}_{n}$, where
    ${\bm{u}}_{i}={\bm{v}}_{\mathrm{pos}(i)},\quad\text{with }\mathrm{pos}(i)=\operatorname{argmax}_{j\in{\mathcal{S}}_{i}}r_{j}.$
    The output ${\bm{u}}_{i}$ is undefined if ${\mathcal{S}}_{i}=\varnothing$.
  • MEAN: The output is a sequence of vectors ${\bm{u}}_{1},\dots,{\bm{u}}_{n}$, where
    ${\bm{u}}_{i}=\operatorname{mean}_{j\in{\mathcal{S}}_{i}}{\bm{v}}_{j}=\frac{1}{|{\mathcal{S}}_{i}|}\sum_{j\in{\mathcal{S}}_{i}}{\bm{v}}_{j}.$
    The output ${\bm{u}}_{i}$ is undefined if ${\mathcal{S}}_{i}=\varnothing$.
  • SINGLE COPY: The output is a sequence of vectors ${\bm{u}}_{1},\dots,{\bm{u}}_{n}$, where
    ${\bm{u}}_{i}={\bm{v}}_{\mathrm{pos}(i)},\quad\text{with }\mathrm{pos}(i)\text{being the unique element in }{\mathcal{S}}_{i}.$
    The output ${\bm{u}}_{i}$ is undefined if $|{\mathcal{S}}_{i}|\neq 1$.
We now impose the following regularity assumption to ensure the feasibility of the operations under consideration:
Assumption C.4 (Regularity Assumption for Attention).
For any input sequence ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, the matrices ${\bm{Q}},{\bm{K}},{\bm{V}}$ and scalars $\rho,\delta$ satisfy the following conditions:
  • For any $i,j\in[n]$, either $|{\bm{q}}_{i}\cdot{\bm{k}}_{j}|\leq\rho$ or ${\bm{q}}_{i}\cdot{\bm{k}}_{j}\leq-\delta$.
  • For any $i,j\in[n]$, either $i=j$ or $|r_{i}-r_{j}|\geq\delta$.
  • The infinity norm of the value matrix ${\bm{V}}$ satisfies $\|{\bm{V}}\|_{\infty}\leq 1$.
Under this assumption, we demonstrate that a logarithmic precision attention layer with $O(d)$ embedding dimension and a single attention head can perform the operations defined in Section C.2.
@@
Suppose Assumption C.4 holds and $\rho\leq\frac{\delta^{2}}{8M}$. For any $\epsilon>0$, there exists an attention layer with a single attention head and $O(d)$ embedding dimension that can approximate the $\mathrm{COPY}$ operation. Furthermore, the $\ell_{\infty}$ norm of the parameters is bounded by $O(\mathrm{poly}(M,1/\delta,\log(n),\log(1/\epsilon)))$.
Formally, for any input sequence ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, let the attention layer outputs be ${\bm{o}}_{1},{\bm{o}}_{2},\dots,{\bm{o}}_{n}$. Then, for any $i\in[n]$ such that ${\mathcal{S}}_{i}\neq\varnothing$, the following holds:
$\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}\leq\epsilon,$
where ${\bm{u}}_{i}$ is the target output of the $\mathrm{COPY}$ operation as defined in Section C.2.
@@
Suppose Assumption C.4 holds and $\rho\leq\frac{\delta\epsilon}{16M\ln(4Mn/\epsilon)}$. For any $0<\epsilon\leq M$, there exists an attention layer with a single attention head and $O(d)$ embedding dimension that can approximate the $\mathrm{MEAN}$ operation. Furthermore, the $\ell_{\infty}$ norm of the parameters is bounded by $O(\mathrm{poly}(M,1/\delta,\log(n),\log(1/\epsilon)))$.
Formally, for any input sequence ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, let the attention layer outputs be ${\bm{o}}_{1},{\bm{o}}_{2},\dots,{\bm{o}}_{n}$. Then, for any $i\in[n]$ such that ${\mathcal{S}}_{i}\neq\varnothing$, the following holds:
$\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}\leq\epsilon,$
where ${\bm{u}}_{i}$ is the target output of the $\mathrm{MEAN}$ operation as defined in Section C.2.
The proofs of Lemmas C.5 and C.6 are omitted here for brevity. Complete proofs can be found in the appendix of Feng et al. (2023).
@@
Suppose Assumption C.4 holds and $\delta-\rho\geq c\rho$ for some constant $c>0$. For any $\epsilon>0$, there exists an attention layer with a single attention head and $O(d)$ embedding dimension that can approximate the $\mathrm{SINGLE~{}COPY}$ operation. Furthermore, the $\ell_{\infty}$ norm of the parameters is bounded by $O(\mathrm{poly}(M,1/\delta,1/c,\log(n),\log(1/\epsilon)))$.
Formally, for any input sequence ${\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{n}$, let the attention layer outputs be ${\bm{o}}_{1},{\bm{o}}_{2},\dots,{\bm{o}}_{n}$. Then, for any $i\in[n]$ such that $|{\mathcal{S}}_{i}|=1$, the following holds:
$\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}\leq\epsilon,$
where ${\bm{u}}_{i}$ is the target output of the $\mathrm{SINGLE~{}COPY}$ operation, as defined in Section C.2.
@@
We construct the query, key, and value vectors as follows:
@@
where $\lambda>0$ is a constant to be determined. Denote $a_{i,j}$ as the attention score, defined as:
$a_{i,j}=\frac{\exp(\lambda({\bm{q}}_{i}\cdot{\bm{k}}_{j}))}{\sum_{j^{\prime}}\exp(\lambda({\bm{q}}_{i}\cdot{\bm{k}}_{j^{\prime}}))}.$
Since $\delta-\rho\geq c\rho$, it follows that $\delta-\rho\geq\frac{c}{c+1}\delta$. Setting
$\lambda=\frac{(c+1)\ln\left(\frac{2nM}{\epsilon}\right)}{c\delta},$
which is bounded by $O(\mathrm{poly}(M,1/\delta,1/c,\log(n),\log(1/\epsilon)))$, we derive the following bounds for $a_{i,\mathrm{pos}(i)}$:
@@
Here, Equation 2 follows from Assumption C.4 and the condition $|{\mathcal{S}}_{i}|=1$, which ensures that for $j^{\prime}\neq\mathrm{pos}(i)$, ${\bm{q}}_{i}\cdot{\bm{k}}_{j^{\prime}}\leq-\delta$; Equation 3 uses the approximation $\frac{1}{1+x}\geq 1-x$ for $x\geq 0$.
Thus, we can bound the error as follows:
$\|{\bm{o}}_{i}-{\bm{u}}_{i}\|_{\infty}$ $=\left\|\sum_{j}a_{ij}{\bm{v}}_{j}-{\bm{v}}_{\mathrm{pos}(i)}\right\|_{\infty}$
$\leq M\|{\bm{V}}\|_{\infty}\cdot\left(1-a_{i,\mathrm{pos}(i)}+\sum_{j\neq\mathrm{pos}(i)}a_{ij}\right)$
$=M\|{\bm{V}}\|_{\infty}\cdot(2-2a_{i,\mathrm{pos}(i)})$
$\leq\epsilon,$
where the last inequality follows from the bound $a_{i,\mathrm{pos}(i)}\geq 1-\frac{\epsilon}{2M}$ and the constraint $\|{\bm{V}}\|_{\infty}\leq 1$ from Assumption C.4. This concludes the proof. ∎
C.3 Technical Lemmas for Constant Precision Calculations
In this section, we establish technical lemmas that underpin constant precision calculations. Assume a system with $2s$-bit fixed-point precision and no exponent bits, and let $B_{s}=2^{s}-2^{-s}$. The largest representable value in this system is $B_{s}$, while the smallest is $-B_{s}$.
Lemma C.8 (Li et al., 2024, Lemmas E.1 and E.2).
For any $s\in\mathbb{N}_{+}$, it holds that $\exp(-B_{s})=0$ and $\exp(B_{s})=B_{s}$.
@@ C.3 Technical Lemmas for Constant Precision Calculations
First, observe that $\exp(B_{s})\geq\mathrm{e}B_{s}>2^{s+1}$. Consequently, $\exp(-B_{s})\leq 2^{-s-1}$, implying $\exp(-B_{s})=0$ due to the truncation to zero under the given precision. For the second claim, note that $\exp(B_{s})\geq B_{s}+1>B_{s}$, which enforces $\exp(B_{s})=B_{s}$ under the constant precision constraints. ∎
@@
For any $s\in\mathbb{N}_{+}$, we have $\mathrm{GeLU}(-B_{s})=0$.
@@
To prove this, it suffices to show that $B_{s}\Phi(-B_{s})\leq 2^{-s-1}$, where $\Phi$ denotes the cumulative distribution function (CDF) of the standard Gaussian distribution.
Case 1: $s=1$. In this case, $B_{s}=\frac{3}{2}$. Thus,
@@
Case 2: $s\geq 2$. For larger $s$, we proceed as follows:
$B_{s}\Phi(-B_{s})$ $=\frac{B_{s}}{\sqrt{2\pi}}\int_{B_{s}}^{+\infty}\mathrm{e}^{-\frac{x^{2}}{2}}\,\mathrm{d}x$
$\leq\frac{B_{s}}{\sqrt{2\pi}}\int_{B_{s}}^{+\infty}\mathrm{e}^{-\frac{B_{s}x}{2}}\,\mathrm{d}x$
$\leq\sqrt{\frac{2}{\pi}}\mathrm{e}^{-\frac{B_{s}^{2}}{2}}\leq\frac{2\sqrt{2}}{\sqrt{\pi}(B_{s}^{2}+2)}$
$\leq\frac{2\sqrt{2}}{\sqrt{\pi}\cdot 2^{2s}}\leq\frac{1}{2^{s+1}}.$
This completes the proof. ∎